3. Suppose that T:V + V and S :V + V such that S is invertible. Prove that T and S-ITS have the same eigenvalues.
Answer #1

3. → Let BB be any basis of V. A= [TB , B = [FS] B. We know that [7]e aud & T has some eigenvalues, Also B & Es} s has same edet ColABA I)= det (B)det (A-XI) det (B) -:- (-: det (59) = det (B)-1) det (A-XI) = 2 BABA and at have same characteristic